Marc Horowitz 21.512

How Metaphorical Structures Relate to Each Other

Lakoff and Johnson claim that metaphor is fundamental to the way people think, speak, and act. They maintain that everything is understood in relationship to something else. In this essay, I will explain how I disagree with one aspect of their argument. I agree that metaphor plays a role in human thought and expression, but I disagree that metaphor structures thought. I believe that thought structures which metaphorical systems are used, specifically, in terms of the Abstraction view which Lakoff and Johnson reject (in Chapter 18).

Take the Argument is War example. Lakoff and Johnson claim that this metaphor expresses how we conceive of argument. But they also claim that arguments and war are different kinds of things. It is with this statement that I disagree. It is because we conceive of argument as a confrontation that the metaphor exists. Argument is verbal confrontation; war is physical confrontation. Because we perceive both concepts as conflicts, we tend to use the same kinds of words and concepts when we talk about them. Lakoff and Johnson claim that we conceptualize arguments in terms of battle. This is a fundamentally different claim than that of conceptualizing battle in terms of argument. I do not see any justification for making one “first” and one “last”. Rather than saying that one form resulted in the other, I would rather say that Argument and War are two different forms of the abstract notion of Conflict, and as such, they draw on the same metaphors.

Another example of Lakoff and Johnson presupposing one system over others is their examples at the end of Chapter 12 (p. 59). They claim that in denotes physical containment, and that when in is used in a social, emotional, or other context, it is metaphorical for the physical meaning. I think one could make an equally valid argument that in denotes social containment, for example, “I am in the class of 1992.” Any physical use is metaphorical. There is no grounds for considering one meaning of in to be basic, and the others to be metaphorical. Here, I would put forth the basic notion of Within. The descriptions “in the kitchen”, “in the Elks”, and “in love” are all just instances of the basic concept of Within. We can use the same metaphors to refer to all instances of Within, and even make up new instances, as long as they are consistent in terms of the way they are viewed.

Lakoff And Johnson present seven reasons why they believe that the Abstraction view is invalid. There are counterarguments to each of their objections.

With respect to the instance of Up-Down orientational metaphors, we can consider each example of an orientational metaphors which we have been given as an instance of a linear scale, with two extremes which can be classified as Good and Bad. Orientational metaphors have meaning because some societal view causes one orientation to be Good and the other Bad. Given this denotation (which is usually arbitrary), we can use words describing different instances of Good and Bad metaphorically. There are directional metaphors, such as “He is in high health, or “the economy is at its peak.” But we might also say “the economy is healthy.” In general, one can exchange any orientational metaphor with any other, and it makes sense. Orientational metaphors are somewhat arbitrary, but as long as we can see which direction a given culture calls good and which it calls bad, we can see the consistency of the metaphorical scheme, even if it is different from our own.

Lakoff and Johnson claim that the Abstraction theory does not distinguish between metaphors of the form A is B and B is A. This is not true. If we take their example of Love is a Journey, I would consider Journey to be a basic concept, of which Love is one instance. There is a ready distinction here, which makes it clear that in our society’s view, Journey metaphors should apply to Love, but the reverse need not and should not be true.

Their third objection is that Abstraction does not explain how a single concept can be described by several different metaphorical systems. They claim that Abstraction requires that a concept have one distinct meta-concept. This is not required. For example, take an abstraction system designed to describe physical objects. A tennis ball is an instance of both Sphere and Yellow. Anything which describes one of these general concepts describes the tennis ball. It is also clear that bananas are not round, even though the tennis ball is yellow. This is absurd, just as using a Container metaphor to describe a war would be absurd.

The fourth objection is that the Abstraction view has no explanation for the tendency to understand the less concrete in terms of the more concrete. This argument comes from a failure to see how abstraction structures work. In their example of Argument is a Building, consider the general concept of Structure. Arguments and Buildings are both kinds of Structures. So, here, we understand Argument through reference to something more concrete which is in the same abstract class as itself. This is similar to the above example, where I might use a tennis ball to demonstrate aspects of the geometry of the earth. One can easily see things about the tennis ball which are not obvious of the earth. The fundamental part of this argument is that although Argument is less concrete than Building, there is an even more general concept which lies above the two.

I do not understand their last three objections. They claim that the abstraction proposal does not have any metaphorical systems, but I have shown that this is not the case. There are metaphorical systems, built up from words which apply directly to certain concepts, and can then be used to describe aspects of related concepts metaphorically. This can be seen in concepts which are related heirarchically. In a Conflict, the participants may attack each other, and so, the word attack can be understandably used in terms of a war or an argument. Because of this, war and argument can be seen as metaphors of each other.

The Abstraction view of metaphor seems to me to be more sensible than the view of Lakoff and Johnson. It is possible to analyze metaphorical systems in terms of abstractions, and there are flaws in all of their arguments agains this view.